Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $40 per unit. Variable expenses are $20.003. Due to an increase in demand, the company estimates that sales will increase by $59,000 during the next year. By how much4. Assume that the operating results for last year were: Sales Variable expenses $ 1,120,000 560,000 Contribution margin Fixe5. Refer to the original data. Assume that the company sold 42,000 units last year. The sales manager is convinced that a 13%b. Would you recommend that the company do as the sales manager suggests? O Yes O No 6. Refer to the original data. Assume ag
Answer #1

Answers

  • Requirement 1 to 5 [total 7]
  • [1]

A

Sale price

$40

B

Variable cost

$20

C = A - B

Contribution margin per unit

$20

D = C/A

CM Ratio

50%

  • [2]

A

Sale price

$40

B

Variable cost

$20

C = A - B

Contribution margin per unit

$20

D = C/A

CM Ratio

50%

E

Fixed cost

180000

F = E/D

Break even point in dollars

$360,000 Answer

  • [3]

Net Operating Income INCREASE by $ 29,500 [which is $ 59000 x 50%]

  • [4a]
    Degree of Operating Leverage = Contribution margin / Net Operating income
    = $ 560000 / 380000
    = 1.47
  • [4b]
    >Increase in sale = 16%
    >Operating Leverage = 1.47
    Net Operating Income increases by = 16 x 1.47 = 23.52%
  • [5a]

Last Year

Proposed

42000

units

63000

units

Total

per unit

Total

per unit

Sales

$1,680,000

$40

$2,192,400

$34.80

Variable cost

$840,000

$20

$1,260,000

$20.00

Contribution margin

$840,000

$20

$932,400

$14.80

Fixed Cost

$180,000

$258,000

Net Income

$660,000

$674,400

  • 5[b]

YES.